# Gas supply shocks, Uncertainty and Price setting: Evidences from Italian Firms

### v3

@@ Abstract
This paper examines how natural gas supply shocks affect Italian firms’ pricing decisions and inflation expectations using quarterly survey data from the Bank of Italy’s Survey on Inflation and Growth Expectations (SIGE). We identify natural gas supply shocks through an external IV-VAR approach exploiting likely unexpected news about interruption to gas supplies to Europe. Our findings show that although gas supply shocks do not have huge effects on gas quantity and only modest effect on gas inventories, they are quickly transmitted to spot electricity prices with persistent effects. We then estimate a proxy internalizing BVAR incorporating firm-level variables from SIGE, documenting that gas supply shocks raise firms’ current and expected prices as well as inflation uncertainty. Finally, we uncover substantial nonlinearities using state-dependent local projections: under high inflation uncertainty, firms successfully pass cost increases on to consumers, sustaining elevated prices; under low uncertainty, recessionary effects dominate, leading firms to cut prices below baseline.
Keywords: Gas supply shocks, Firms’ Price Setting, Inflation Uncertainty, Firms Expectations, Local Projections, Bayesian Vector Autoregression, Instrumental Variables
@@ 1 Introduction
Unlike the extensive literature on oil supply shocks, research examining the role of natural gas supply disruptions has emerged only recently. In this paper, we aim at identifying a shock to the supply of natural gas and at evaluating the impact of this shock on firms price setting behavior. We identify the shock following an external IV-VAR approach following kanzig2021macroeconomic. We construct the instrument by exploiting episodes that carried likely exogenous news about supply distruptions affecting Europe, and in particular Italy. Our identification strategy is hence close in the spirit to the one followed by alessandri2025natural, which instead adopt an internal instrument and consider a larger number of surprises. Then, we evaluate the effect of this shock on firms’ pricing behavior and expectations using the generalized impulse response functions from a proxy internalizing VAR, using the shock identified in the first VAR as an external instrument. Additionally, we evaluate non linear effects in the transmission of the shock to firms’ prices by using state dependent local projections framework (ramey2018government; F2021). Our contribution to the literature lies in examining how natural gas supply shocks affect firm-level outcomes and the pass-through using firm level micro data. We base our analysis on quarterly data from the Bank of Italy’s Survey on Inflation and Growth Expectations (SIGE).
@@
We draw from a vast literature about macroeconomic outcomes of energy related shocks. A large number of papers study the macroeconomic effects of oil shocks (hamilton1983oil; kilian2009not; caldara2019oil; conflitti2019oil; kilian2022oil). Our work is closely related to the recent papers which study the effect of gas supply shocks, focusing on the recent inflationary surge in European economies. alessandri2025natural identify a natural gas supply shock using daily news on European gas market as an instrument. boeck2025natural and lopez2025pass estimate the pass through to inflation of gas price shocks. adolfsen2024gas use a BVAR framework to identify shocks driving the natural gas market in the EU. They document that the pass through is heterogeneous depending on the shock type. guntner2024sudden study the effect of recent supply disruptions in natural gas supply in the German market. casoli2024energy study the interaction between oil/gas shocks and their effects on inflation in the Euro area. With respect to those authors, we focus on firm level expectations and price setting behavior.
@@ 2 Data: Survey on Inflation and Growth Expectations
This section illustrates the features of the Italian Survey on Inflation and Growth expectations (SIGE) [^1] Survey on Inflation and Growth Expectations, which constitutes our primary data source for firm expectations and price-setting decisions. As one of the longest-running firm-level expectation surveys in a G7 country, SIGE provides a rich source of information on both expectations formation and pricing behavior.
The survey is conducted by Bank of Italy at a quarterly frequency starting from the end of 1999. It provides a rotating panel collecting different kind of information from Italian firms. Among others, firms are asked to provide a point estimate for their year on year inflation expectations, and the expected change in their own prices over the course of the year. Furthermore, firms report the average change in their own realized price over the year, allowing a comparison between expectations and outcomes. Additionally, firms respond to a range of categorical question related to what are the main factors that will affect their own prices over the course of the next year. In this case, responses are on a scale giving information about both direction (downward or upward pressure) and intensity (ranging from strong to modest). In the following paragraphs, we describe the aggregate information that we extract from the survey in order to be used in our analysis.
@@
The observed patterns suggest the presence of a cost-push narrative during the European energy crisis that followed the post-pandemic recovery: raw material and intermediate input costs are cited as primary pricing drivers starting in 2021, followed by labor costs and demand pressures. Notably, inflation expectations themselves become a significant pricing factor only around 2022, suggesting that as firms observed widespread price increases, expectations began feeding directly into their own pricing decisions. Given the prominence of input costs—particularly energy-related raw materials—as key drivers of pricing behavior during this period, this motivates our focus on examining the effects of natural gas supply shocks on firm-level outcomes.
@@
where $f_{l}$ is assumed to be normal, $n_{h}$ is the number of firms which responses belong to $S_{h}$, $N_{c}$ is the number of non rounded responses and $\theta=(\phi_{h},\lambda_{t},\mu_{l},\sigma_{l})$ is a vector of parameters . The mixture weight $\lambda_{t}$ is the portion of uncertain respondents.555binder2017measuring includes non respondents and measures the proportion of uncertain households as $U_{t}=\frac{\lambda_{t}N_{t}+\gamma N_{t}^{DK}}{N_{t}+\gamma N_{t}^{DK}}$ where $N_{t}$ is the total number of respondents, $N_{t}^{DK}$ is the number of respondents which answer is ”don’t know” and $\gamma$ is a parameter which measures their weight. As in the SIGE survey the question about expected inflation is mandatory, we do not have DKtypeDK-type respondents in our sample and hence we measure the proportion of uncertain firms by simply taking $\lambda_{t}$. The vector of parameters $\theta_{t}$ is estimated via maximum likelihood. Figure 4 reports the inflation uncertainty index between the last quarter of 1999 and the second quarter of 2025. Notably, since firms are informed about current price levels at the time of the interview, our index isolates doubts about future inflation, abstracting from any confusion about the current state of prices.
@@
Inflation uncertainty among informed firms is notably lower than that typically observed for consumers (binder2017measuring). During the early sample period, uncertainty remains low, reflecting stable price dynamics. The first major spike occurs in 2008Q3 during the Global Financial Crisis, followed by persistently high levels throughout the European sovereign debt crisis. Subsequent surges are observed around the Russian invasion of Crimea (2014Q3), Greece’s IMF default (2015Q2), and during the Italian banking crisis (2016Q3). The fall of the Conte II government in 2020Q4 coincides with a sharp rise in uncertainty, likely reflecting concerns over Italy’s access to European recovery funds amid the COVID-19 pandemic. A major increase appears in 2022Q1 with Russia’s invasion of Ukraine, driven by severe energy supply disruptions and their inflationary effects. Uncertainty rises again in 2022Q4, as Italian inflation reaches its highest level since the adoption of the euro, underscoring widespread concerns about the persistence and spread of price pressures.
@@ 3.1 Identification through external instrument
In order to identify a supply shock to natural gas, we follow the approach of kanzig2023unequal (kanzig2021macroeconomic, kanzig2023unequal). As alessandri2025natural, we exploit news extracted from Refinitiv. Differently from them, we focus on a smaller number of news from Reuters about interruptions of gas supplies to Europe which are highly likely to be exogenous to the economy and unexpected by the markets. We include episodes of unexpected interruptions caused by extreme weather event, like the cold snap that hit Russia in 2012 forcing it to reduce the export of natural gas to satisfy its domestic demand, and interruptions in the supply caused by terrorist attacks, protests and unplanned maintenance in Libya and Algeria. Furthermore, we include incidents like the explosion that hit the Baumgarten natural gas facility in Austria in 2017 causing the interruption of the flow of gas to the Italian Hub of Tarvisio and the fire in the Freeport liquefied natural gas facility in US which reduced the supply of LNG to Europe. Additionally, for the 2022 period we include some announcements from Russia which have been perceived as a surprise from the market and which contain news about future reduction in the supply of natural gas. As an example, the announcement by Putin that Russia was going to stop the deliveries of natural gas, unless the payments were made in robles in March 2022 concretely reduced the amount of purchasable commodity for countries which could not or did not want to comply with the request, and has been followed by a positive movement of the Dutch TTF (alessandri2025natural). We avoid to include the sabotage of the Nord Stream 2 pipeline since the latter never actually delivered natural gas to Europe, thus the event did not concretely change the amount of commodity supplied to European countries. We also exclude the strikes in gas production facilities in Norway during 2022, since these episodes were motivated by the request of Norvegian workers to increase wages in response to inflationary pressures. Furthermore, to reduce the contamination of the instrument with geopolitical shocks we do not include the Russian invasion of Ukraine in February 2024 even though it has been followed by a massive surge in the Dutch TTF.

### v4

# Gas supply shocks, Uncertainty and Price setting: Evidence from Italian Firms

@@ Abstract
This paper examines how natural gas supply shocks affect Italian firms’ pricing decisions and inflation expectations using quarterly survey data from the Bank of Italy’s Survey on Inflation and Growth Expectations (SIGE). Our findings show that gas supply shocks are quickly transmitted to spot electricity prices with persistent effects. We then estimate a proxy internalizing BVAR incorporating firm-level variables from SIGE, documenting that gas supply shocks raise firms’ current and expected prices as well as inflation uncertainty. Additionally, we document heterogeneous effects across sectors. Finally, we uncover substantial nonlinearities using state-dependent local projections: under high inflation uncertainty, firms successfully pass cost increases on to consumers, sustaining elevated prices; under low uncertainty, recessionary effects dominate, leading firms to cut prices below baseline.
Keywords: Gas supply shocks, Price Setting, Inflation Uncertainty, Firms Expectations, Local Projections, Bayesian Vector Autoregression, Instrumental Variables
@@ 1 Introduction
Unlike the extensive literature on oil supply shocks, research examining the role of natural gas supply disruptions has emerged only recently. In this paper, we aim at identifying a shock to the supply of natural gas and at evaluating the impact of this shock on firms price setting behavior. We identify the shock following an external IV-VAR approach following kanzig2021macroeconomic. We construct the instrument by exploiting episodes that carried likely exogenous news about supply disruptions affecting Europe, and in particular Italy. Our identification strategy is hence similar to the one followed by alessandri2025natural and by colombo2025understanding. Then, we evaluate the effect of this shock on firms’ pricing behavior and expectations by simulating impulse response functions from a recursive proxy internalizing VAR, using the shock identified in the first VAR as an internal instrument. We focus on firms belonging to the manufacturing and services sector, documenting important heterogeneity in the pass-trough of the shock to firms’ prices. Additionally, we evaluate the presence of nonlinearities in the pass-trough by using state dependent local projections framework (ramey2018government; F2021).
@@
We draw from a vast literature about macroeconomic outcomes of energy related shocks. A large number of papers study the macroeconomic effects of oil shocks (hamilton1983oil; kilian2009not; caldara2019oil; conflitti2019oil; kilian2022oil). Our work is closely related to the recent papers which study the effect of gas supply shocks, focusing on the recent inflationary surge in European economies. alessandri2025natural identify a natural gas supply shock using daily news on European gas market as an instrument. colombo2025understanding use a similar approach for identifying gas supply shocks and extend it by identifying gas demand shocks for the US and for the Euro Area by exploiting anomalies in temperature variations. boeck2025natural identify a gas price shock for the Euro Area by using a combination of sign and zero restrictions, while lopez2025pass use sign and narrative restriction. Both of them estimate the pass through of gas price shocks to inflation. adolfsen2024gas use a BVAR with sign and narrative restrictions to identify shocks driving the natural gas market in the EU. They document that the pass through is heterogeneous depending on type of gas market related shocks hitting the economy. guntner2024sudden study the effect of recent supply disruptions in natural gas supply in the German market. casoli2024energy study the interaction between oil/gas shocks and their effects on inflation in the Euro area. With respect to those authors, we focus on Italy and on firm level expectations and price setting behavior.
@@ 2 Data: Survey on Inflation and Growth Expectations
This section illustrates the features of the Italian Survey on Inflation and Growth expectations (SIGE) [^1] https://www.bancaditalia.it/pubblicazioni/indagine-inflazione/, which constitutes our primary data source for firm expectations and price-setting decisions. As one of the longest-running firm-level expectation surveys in a G7 country, SIGE provides a rich source of information on both expectations formation and pricing behavior.
The survey is conducted by the Bank of Italy at a quarterly frequency starting from the end of 1999. It is a rotating panel that collects different kinds of information from Italian firms with at least 50 employees. The survey is stratified by (i) number of employees, (ii) geographical area, and (iii) sector. Among others, firms are asked to provide a point estimate for their year on year inflation expectations, and the expected change in their own prices over the course of the year. Furthermore, firms report the average change in their own realized price over the year, allowing a comparison between expectations and outcomes. Additionally, firms respond to a range of categorical question related to what are the main factors that will affect their own prices over the course of the next year. In this case, responses are on a scale giving information about both direction (downward or upward pressure) and intensity (ranging from strong to modest). In the following paragraphs, we describe the aggregate information that we extract from the survey in order to be used in our analysis.
@@
The evidence points to a clear cost-push narrative during the European energy crisis. The index for raw-material prices displays a consistently positive contribution over the entire sample, but it surges during the energy crisis, almost reaching the upper bound of the scale. This indicates an almost universal perception among firms that rising input costs were exerting strong upward pressure on their selling prices. Inflation expectations became the second most relevant upward pricing factor only around 2021, when the importance of intermediate-input costs also intensified. This pattern suggests that as firms observed widespread price increases, expectations began to feed directly into their own pricing decisions. Given the prominence of input costs—particularly energy-related raw materials—as key drivers of inflation during the post pandemic recovery, this motivates our focus on examining the effects of natural gas supply shocks on firm-level outcomes.
@@
(a) Price changes
(b) Expected price changes
Figure 4: Price realizations and expectations across sectors. The figure plots average price changes over the last 12 months (left) and the average expected price changes over the next 12 months (right) for Italian firms in the manufacturing (red dotted line) and services sectors (lightblue dashed line), compared with the overall series (blue solid line). Source: Bank of Italy, Survey on Inflation and Growth Expectations. Sample: 1999:Q4–2025:Q2.
In Figure 4, we compare price expectations and realizations for firms in the manufacturing and services sectors. The figures reveal a more volatile and reactive pricing behavior in manufacturing, whereas services adjust more gradually and with a slower pass-through of shocks. Until 2020, expected and realized price changes evolve similarly across sectors, following comparable cyclical patterns. The divergence emerges with the energy crisis: manufacturing firms respond much more forcefully, with both expected and realized prices rising sharply and reaching substantially higher levels than in services. This suggests that manufacturing incorporates shocks more rapidly and with greater intensity. After 2022, the adjustment in manufacturing is also more abrupt, while the correction in services is smoother and more persistent. In the post-crisis period, realized price changes in services remain comparatively higher, indicating greater rigidity than in manufacturing.
Price setting margins
We exploit firms’ answers on how much their prices changed over the previous 12 months to construct proxies of the extensive and intensive margins of price adjustment. The extensive margin is defined as the share of firms reporting a non–negligible year-on-year change in prices, which provides a survey–based measure of the incidence of price adjustment. The intensive margin is the average reported year-on-year price change among firms that adjusted. These indicators do not track the precise timing or the number of price changes within the year, but they summarize whether firms adjusted at least once over the past 12 months and by how much, and therefore offer meaningful proxies for the underlying price-setting margins.
Let $\Delta p_{i,t}^{yoy}$ denote the year-on-year percentage change in the price reported by firm $i$ in quarter $t$. We define our proxy of the extensive margin $F_{t}$ as the share of firms reporting a non-negligible price change:
$F_{t}=\frac{1}{N_{t}}\sum_{i=1}^{N_{t}}\mathbf{1}\!\left(|\Delta p_{i,t}^{yoy}|\geq\varepsilon\right)$
where $\varepsilon$ is a small threshold used to filter out reporting noise. The intensive margin $I_{t}$ is instead defined as the average reported year-on-year price change among adjusting firms:
$I_{t}=\mathbb{E}\!\left[\Delta p_{i,t}^{yoy}\,\middle|\,|\Delta p_{i,t}^{yoy}|\geq\varepsilon\right].$
Under this definition, the unconditional mean of firms’ year-on-year price changes satisfies:
$\mathbb{E}[\Delta p^{yoy}_{i,t}]=F_{t}\,I_{t},$
The series, represented by the solid blue lines in Figure 5, exhibit cyclical patterns that are consistent with the micro-pricing evidence and provide a characterization of the price adjustment behavior of Italian firms over time. For both of them, we observe positive comovements with realized inflation. Before 2008, roughly 80% of firms in the sample adjusted their prices at least once per year, and the magnitude of these adjustments remained moderate. In 2008, both margins collapse: fewer firms change their prices, and those that do implement substantial price cuts, marking a sharp and broad-based disinflationary episode aligned with the onset of the global financial crisis. Following this break, the period from 2011 to 2019 is characterized by persistently muted pricing behavior. The extensive margin settles at historically low levels, while the intensive margin hovers around zero, indicating that price adjustments—when they occur—are infrequent, small, and lack a clear direction. This pattern reflects a prolonged environment of weak inflation and subdued demand. After 2020, the dynamics shift abruptly. The extensive margin rises markedly, signaling that a much larger share of firms resumes active price adjustment, while the intensive margin reaches unprecedented highs, pointing to large and predominantly upward price changes. These developments reveal a clear regime shift towards strong and widespread price increases driven by the post-pandemic recovery, the surge in energy costs, and the European gas supply shock.
(a) Extensive margin
(b) Intensive margin
Figure 5: SIGE intensive and extensive margins The figure plots the proxies for the extensive (left) and extensive (right) margins from the SIGE survey responses about the reported price changes in the last 12 months for Italian firms. Source: Bank of Italy, Survey on inflation and growth expectations. Sample: 1999:Q4–2025:Q2.
Computing the price-setting margins by differentiating across sectors reveals clear patterns. First, compared with the services sector, a slightly higher share of manufacturing firms adjust their prices over the sample. The gap between the two groups widens from 2013 until the onset of the pandemic. Manufacturing firms also start increasing their prices more frequently after 2020. In 2022, almost 90% of manufacturing firms adjusted their prices. By contrast, in the services sector the increase in the intensive margin is more gradual and contained, although still substantial.
The intensive margin is similar for both sectors for most of the sample, but after 2020 clear differences emerge: manufacturing firms adjust their prices more aggressively until 2023, after which their intensive margin decreases sharply. In contrast, price adjustments by firms in the services sector remain more contained overall, but the size of their adjustments surpasses that of manufacturing firms after 2023.
@@
where $f_{l}$ is assumed to be normal, $n_{h}$ is the number of firms which responses belong to $S_{h}$, $N_{c}$ is the number of non rounded responses and $\theta=(\phi_{h},\lambda_{t},\mu_{l},\sigma_{l})$ is a vector of parameters . The mixture weight $\lambda_{t}$ is the portion of uncertain respondents.555binder2017measuring includes non respondents and measures the proportion of uncertain households as $U_{t}=\frac{\lambda_{t}N_{t}+\gamma N_{t}^{DK}}{N_{t}+\gamma N_{t}^{DK}}$ where $N_{t}$ is the total number of respondents, $N_{t}^{DK}$ is the number of respondents which answer is ”don’t know” and $\gamma$ is a parameter which measures their weight. As in the SIGE survey the question about expected inflation is mandatory, we do not have DKtypeDK-type respondents in our sample and hence we measure the proportion of uncertain firms by simply taking $\lambda_{t}$. The vector of parameters $\theta_{t}$ is estimated via maximum likelihood.
@@
Figure 6 reports the inflation uncertainty index between the last quarter of 1999 and the second quarter of 2025. Notably, since firms are informed about current price levels at the time of the interview, our index isolates doubts about future inflation, abstracting from any confusion about the current state of prices. [^6] A concern may be that building the uncertainty index only on firms who receive the information about the current level of prices may lead to an under measurement in uncertainty. In appendix A we show that the share of uncertain firms who do not receive information about current inflation when interviewed is similar both in size and temporal patterns to our baseline. Inflation uncertainty among informed firms is notably lower than that typically observed for consumers (binder2017measuring). During the first part of the sample, uncertainty remains low, reflecting stable price dynamics. The first major spike occurs in 2008Q3 during the Global Financial Crisis, followed by persistently high levels throughout the European sovereign debt crisis. Subsequent surges are observed around the Russian invasion of Crimea (2014Q3), Greece’s IMF default (2015Q2), and during the Italian banking crisis (2016Q3). The political crisis that led to the fall of the Conte II government in 2020Q4 coincides with a sharp rise in uncertainty, likely reflecting concerns over Italy’s access to European recovery funds amid the COVID-19 pandemic. A major increase appears in 2022Q1 with Russia’s invasion of Ukraine, driven by severe energy supply disruptions and their inflationary effects. Uncertainty rises again in 2022Q4, as Italian inflation reaches its highest level since the adoption of the euro, underscoring widespread concerns about the persistence and spread of price pressures. Figure 7 compares our inflation-uncertainty index with several series of interest.
Figure 7: SIGE Inflation Uncertainty Index and related series. The figure plots the inflation uncertainty index for firms computed by adopting the framework of binder2017measuring against the standard deviation of firms’ inflation expectations from SIGE (top-left), a measure of firms’ inflation uncertainty obtained by applying the procedure of Jurado2015 to the first moment of firms’ inflation expectations from SIGE and exploiting the high dimensional dataset provided by barigozzi2024large (top-right), the annualized HICP for Italy (center-left), the Economic Policy uncertainty index of baker2016measuring (center-right) and the volatility of realized inflation, estimated by adopting an ARIMA-GARCH model on the series of the annualized HICP for Italy. All series are standardized for comparability. Source: Bank of Italy, Survey on inflation and growth expectations, EUROSTAT, baker2016measuring, barigozzi2024large. Sample: 1999:Q4–2025:Q2.
First, we consider two alternative indicators of firms uncertainty about future inflation: the cross-sectional standard deviation of inflation expectations and an uncertainty index constructed following the methodology of Jurado2015, by exploiting the high-dimensional dataset of macroeconomic time series for the Italian economy provided by barigozzi2024large. [^7] No outlier correction is performed, and missing data are imputed using the algorithm of mccracken2016fred. These two measures provide natural benchmarks, as they capture either disagreement among firms or the volatility of the unpredictable component of macroeconomic dynamics. Our measure is closely aligned with these firm-specific indicators. Its correlation with the standard deviation of inflation expectations is 85.6%, and the correlation with the Jurado2015 style index is 70.1%. By contrast, correlation with annual inflation is 59.5%, and the correlation with the Economic Policy Uncertainty index of baker2016measuring is only 13.8%. This suggests that our measure is not simply reflecting changes in the level of inflation, nor is it driven by broader political or news-based uncertainty. Instead, it captures a distinct dimension of uncertainty that is specific to firms’ inflation expectations. Additionally, we compare our index with an estimate of realized inflation volatility estimated by using an ARIMA-GARCH model on annualized inflation for Italy. In this latter case, correlation with our index is 75%.
@@ 3.1 Identification through external instrument
In order to identify a supply shock to natural gas, we follow the approach of kanzig2023unequal (kanzig2021macroeconomic, kanzig2023unequal) based on the combination of narrative surprises and high frequency identification. As alessandri2025natural, we exploit news extracted from Refinitiv. Differently from them, we focus on a small number of news from Reuters about interruptions of gas supplies to Europe, and especially to Italy, which are highly likely to be exogenous to the economy and unexpected by the markets. Hence, we are closer to colombo2025understanding in the spirit. We include episodes of unanticipated interruptions caused by terrorist attacks, protests and unplanned maintenance in Libya and Algeria, as these two countries are important suppliers of natural gas for Italy. Furthermore, we include incidents like the explosion that hit the Baumgarten natural gas facility in Austria in 2017 causing the interruption of the flow of gas to the Italian natural gas hub in Tarvisio and the fire in the Freeport liquefied natural gas facility in US which reduced the supply of LNG to Europe. Additionally, for the 2022 period we include some announcements from Russia which have been perceived as a surprise from the market and which contain news about reduction in the supply of natural gas. As an example, the announcement by Putin that Russia was going to stop the deliveries of natural gas, unless the payments were made in robles in March 2022 concretely reduced the amount of purchasable commodity for countries which could not or did not want to comply with the request, and has been followed by a positive movement of the Dutch TTF (alessandri2025natural). We avoid to include the sabotage of the Nord Stream 2 pipeline since the latter never actually delivered natural gas to Europe, thus the event did not concretely change the amount of commodity supplied to European countries. For the same reason, we do not include the stop of gas supplies to Ukraine in 2014, as they did not affect deliveries to Europe (lopez2025pass). We also exclude the strikes in gas production facilities in Norway during 2022, since these episodes were motivated by the request of Norvegian workers to increase wages in response to inflationary pressures. Furthermore, to reduce the contamination of the instrument with geopolitical shocks we do not include the Russian invasion of Ukraine in February 2022 even though it has been followed by a massive surge in the Dutch TTF.
@@ 14 June 2022 NS1 operate at limited capacity due to turbines stuck in Canada
19 August 2022 Gazprom announces that Nord Stream 1 three days shut down for maintenance on the remaining operational turbine